OpenStudy (paradise):

Two cards are drawn with replacement from a standard deck of 52 playing cards. Find the probability that the first card is a king and the second is a diamond.

OpenStudy (amistre64):

how many kings are in a deck? how many diamonds are in a deck?

OpenStudy (paradise):

Its doesn't say

OpenStudy (amistre64):

these questions assume you know how a deck of cards looks like

OpenStudy (amistre64):

normally: there are 4 suits in a deck of cards, hearts, diamonds, spades, and clubs 52/4 = 13 cards in each suit, so 13 diamonds out of 52 cards each suit has numbered cards and face cards: ace, 2, 3, 4, 5, 6, 7, 8, 9, jack, queen, king since there are 4 suits, and each suit has a king; there are 4 kings out of 52 cards

OpenStudy (amistre64):

the next question is: how do you solve a probability question that asks for an "AND" relationship

OpenStudy (paradise):

Would you have to add them together?

OpenStudy (paradise):

(4/52)(13/52) + (13/52)(4/52) like that?

OpenStudy (amistre64):

thats not quite correct, but a good effort. the textbooks give us:\[P(AnB)=P(A)*P(B)\] so that would be multiply them together

OpenStudy (amistre64):

\[P(king)=\frac{4}{52}=\frac1{13}\] \[P(diamond)=\frac{13}{52}=\frac1{4}\] \[P(knd)=\frac1{13}*\frac14\]